OpenStudy (anonymous):

complete the square both numbers x^(2)+(3)/(4)x+_?_=(x+_?_)^(2)

myininaya (myininaya):

i already did this

OpenStudy (anonymous):

it wasnt right it said

OpenStudy (anonymous):

i tried it

myininaya (myininaya):

\[x^2+\frac{3}{4}x+(\frac{3}{4 \cdot 2})^2=(x+\frac{3}{4 \cdot 2})^2\]

myininaya (myininaya):

you have to simplify i didn't do that for you

OpenStudy (mertsj):

For crying out loud. I did it for him too.

myininaya (myininaya):

\[(\frac{3}{4 \cdot 2})^2=(\frac{3}{8})^2=\frac{3^2}{8^2}=\frac{9}{64}\]

OpenStudy (anonymous):

thats it simplified aw man no wonder i dont think thats all though

myininaya (myininaya):

\[x^2+\frac{3}{4}x+(\frac{3}{4 \cdot 2})^2=(x+\frac{3}{4 \cdot 2})^2 \] this is all \[x^2+\frac{3}{4}x+\frac{9}{64}=(x+\frac{3}{8})^2\]

OpenStudy (mertsj):

what do you use to write fractions and equations the way you do Myininaya?

OpenStudy (anonymous):

thnk you so very much :)

myininaya (myininaya):

do you see the equation thingy below the box you are typing in i use that sometimes frac{ a}{b} will give \[\frac{a}{b}\]

OpenStudy (anonymous):

magic. she uses magic

OpenStudy (mertsj):

ty

OpenStudy (anonymous):

if you want to see the code, right click over the expression and choose "show source"

OpenStudy (mertsj):

very cool

OpenStudy (anonymous):

\[\color{#FF0099}{\overrightarrow{\text{myininaya}}}\]

OpenStudy (anonymous):

that is a better color

OpenStudy (anonymous):

i think you can use all web colors

OpenStudy (anonymous):

you can pick your own. it is nice. now i have a question. will we ever reach level 101?

myininaya (myininaya):

that is a question for shadowy